OpenStudy (anonymous):

given f(x) = 1/x, find f(1+Δx)-f(1) / (Δx)

OpenStudy (anonymous):

i just need help on how to write the equation!

OpenStudy (anonymous):

so i plug in 1/x inside of ever f( )

OpenStudy (anonymous):

does that mean for the numerator i get ((1+Δx)/1) - (1/x) / Δx?

OpenStudy (helder_edwin):

just replace \[ \large \frac{f(1+\Delta x)-f(1)}{\Delta x}= \frac{\frac{1}{1+\Delta x}-\frac{1}{1}}{\Delta x}= \]

OpenStudy (anonymous):

hmmm. seems i did it wrong

OpenStudy (anonymous):

we can call \(\Delta x\) \(h\) to make typing simpler and so you are looking for \[\frac{\frac{1}{1+h}-\frac{1}{1}}{h}\]

OpenStudy (helder_edwin):

no. it is the other way around: plug \(1+\Delta x\) into f(x)

OpenStudy (anonymous):

your main work is this: subtracting \[\frac{1}{1+h}-1\] that is the only hard part

OpenStudy (helder_edwin):

yes. satellite73 is right.

OpenStudy (anonymous):

you need \(1+h\) in the denominator, so compute via \[\frac{1-(1+h)}{1+h}\] distribute carefully to get \[\frac{-h}{1+h}\]

OpenStudy (anonymous):

then (and only then) recall that you are dividing all this by \(h\) which gives \[\frac{-1}{1+h}\]